This is a quadratic word problem.
Find two consecutive even intergers such that the square of the smaller is 10 more than the larger.

Answer by ankor@dixie-net.com(22740)

Find two consecutive even integers
x, (x+2)
:
such that the square of the smaller is 10 more than the larger
x^2 = (x+2) + 10
x^2 = x + 12
Arrange as a quadratic equation
x^2 - x - 12 = 0
Factor
(x-4)(x+3) = 0
Two solutions
x = 4; then the next even integer = 6; (16 = 6 + 10)
:
x = -3 Not a solution, it's an odd number